OpenStudy (anonymous):

I am confused with dividing square roots, I have a problem with it and I want to know if there is anyone can clear it up and show me how to divide square roots.

OpenStudy (anonymous):

Can someone explain how to divide square roots?

OpenStudy (anonymous):

Generally you don't want square roots in your equation, but if you do have square roots the you can use the following property: \[\sqrt{2}=2^{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

I don't get it, =\

OpenStudy (anonymous):

can you explain it better

OpenStudy (anonymous):

because i am still in confusion.

OpenStudy (jhannybean):

\[\huge \sqrt[\color{red}{2}]{9}\]You see the small little red 2 ontop of the square root?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ \sqrt{2} }=\sqrt{\frac{ 2 }{ 2 }}=\sqrt{1}\]

OpenStudy (anonymous):

you can also use this property

OpenStudy (jhannybean):

It tells you not only how many of the similar pairs you can extract from the square root, it also tells you to what power the number inside is at.

OpenStudy (jhannybean):

so think of \(9\) as a value of x. (x usually represents a number value). m is any power of the number inside, and # is just like that little red two from the previous post. We can state that \[\huge \sqrt[\#]{x^m} = (x)^{m/ \#}\]

OpenStudy (jhannybean):

This is the general format of writing square roots as fractional powers.

OpenStudy (jhannybean):

Are you somewhat understanding it?

OpenStudy (anonymous):

somewhat

OpenStudy (jhannybean):

So now we can take our \(\sqrt{9}\). \[\huge \sqrt[2]{9^1} = (9)^{1/2}\] Do you see how this works?

OpenStudy (jhannybean):

Just like the general format we stated above!

OpenStudy (anonymous):

oh ok, yeah get it better, thank you

OpenStudy (jhannybean):

Yeah? :) So let's see an example!

OpenStudy (jhannybean):

You had one posted earlier but you removed it :(

OpenStudy (anonymous):

umm.. example, i am not ready for an example though =3

OpenStudy (anonymous):

i would post the example but am not ready for one

OpenStudy (jhannybean):

Ahh, ok. then I will give you one! you tell me how to write it as a fraction, and write the fraction ones as square roots!

OpenStudy (anonymous):

im not sure i can do it

OpenStudy (anonymous):

=(

OpenStudy (jhannybean):



OpenStudy (jhannybean):

So first question you ask yourself, "what is the power ontop of the square root sign?"

OpenStudy (jhannybean):

We just talked about this! I believe you can do it :)

OpenStudy (anonymous):

3 to 5th power, would that be how im supposed to answer??

OpenStudy (anonymous):

struggling

OpenStudy (jhannybean):

no no, what is the power ontop of the square root?

OpenStudy (jhannybean):

\(3^5\) is partially right, you are almost there :) We have to take it one step at a time!

OpenStudy (anonymous):

?the 5th power?

OpenStudy (anonymous):

or something

OpenStudy (anonymous):

nice explanation jhanny!

OpenStudy (jhannybean):

Remember, ALWAYS REMEMBER THIS. the base POWER of a square root is ALWAYS ALWAYS ALWAYS 2!

OpenStudy (jhannybean):